How many moles of magnesium oxide are produced by the reaction of 3.82 g of magnesium nitride with 7.73 g of water? The reaction is: Mg3N2(s) + 3H20(g) → 2N Halg) + 3MgO(s) 0.113 mol 0.0378 mol 0.429 mo 0.0756 mol 4.57 mol Question5 1 pts Solid carbon reacts with chlorine gas to give gaseous carbon tetrachloride. C (s) + 2 Cl2 (g) → CCl4 (g) If 150 g of carbon reacts with 250 g Cl2 with an 85% yield, what is the mass of CCI4 produced? O 182.7 8 230.5 g 274.4 8
Answer #1

1)

Molar mass of Mg3N2 = 3*MM(Mg) + 2*MM(N)

= 3*24.31 + 2*14.01

= 100.95 g/mol

mass of Mg3N2 = 3.82 g

we have below equation to be used:

number of mol of Mg3N2,

n = mass of Mg3N2/molar mass of Mg3N2

=(3.82 g)/(100.95 g/mol)

= 3.784*10^-2 mol

Molar mass of H2O = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass of H2O = 7.73 g

we have below equation to be used:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(7.73 g)/(18.016 g/mol)

= 0.4291 mol

we have the Balanced chemical equation as:

Mg3N2 + 3 H2O ---> 3 MgO + 2 NH3

1 mol of Mg3N2 reacts with 3 mol of H2O

for 3.784*10^-2 mol of Mg3N2, 0.1135 mol of H2O is required

But we have 0.4291 mol of H2O

so, Mg3N2 is limiting reagent

we will use Mg3N2 in further calculation

Molar mass of MgO = 1*MM(Mg) + 1*MM(O)

= 1*24.31 + 1*16.0

= 40.31 g/mol

From balanced chemical reaction, we see that

when 1 mol of Mg3N2 reacts, 3 mol of MgO is formed

mol of MgO formed = (3/1)* moles of Mg3N2

= (3/1)*3.784*10^-2

= 0.113 mol

Answer: 0.113 mol

Only 1 question at a time please